# For Better or For Worse? Learning Minimum Variance Features With Label Augmentation

### v1

@@ Abstract
Data augmentation has been pivotal in successfully training deep learning models on classification tasks over the past decade. An important subclass of data augmentation techniques - which includes both label smoothing and Mixup - involves modifying not only the input data but also the input label during model training. In this work, we analyze the role played by the label augmentation aspect of such methods. We prove that linear models on linearly separable data trained with label augmentation learn only the minimum variance features in the data, while standard training (which includes weight decay) can learn higher variance features. An important consequence of our results is negative: label smoothing and Mixup can be less robust to adversarial perturbations of the training data when compared to standard training. We verify that our theory reflects practice via a range of experiments on synthetic data and image classification benchmarks.
@@ 1 Introduction
Perhaps the two most widely applied data augmentation methods in this subcategory are label smoothing (Szegedy et al., 2015) and Mixup (Zhang et al., 2018). Label smoothing replaces the one-hot encoded labels in the training data with smoothed out labels that assign non-zero probability to every possible class (see Section 2 for a formal definition). Mixup similarly smooths out the training labels, but does so via introducing random convex combinations of data points and their labels. As a result, Mixup modifies not only the training labels but also the training inputs. The general principles of label smoothing and Mixup have been extended to create several variants, such as Structural Label Smoothing (Li et al., 2020), Adaptive Label Smoothing (Wang et al., 2021), Manifold Mixup (Verma et al., 2019), CutMix (Yun et al., 2019), PuzzleMix (Kim et al., 2020), SaliencyMix (Uddin et al., 2020), AutoMix (Liu et al., 2021), and Noisy Feature Mixup (Lim et al., 2022).
@@
1.1 Outline of Main Contributions
Our main contribution can be summarized as,
For linearly separable data distributions that can be learned using any one of multiple features, models trained with Mixup or label smoothing learn only the minimum variance feature.
In order to work up to this claim, we provide the necessary background on Mixup and label smoothing and formalize the notion of data distributions with multiple separating features in Section 2. Our definition draws heavily from the idea of multi-view data introduced by Allen-Zhu & Li (2021), but is considerably simpler as it does not involve patches or feature noise.
We then prove our main results in Section 3, where we show that linear models trained with Mixup or label smoothing on the aforementioned type of binary class data distribution are correlated only with the minimum variance feature in the data. On the other hand, we also show that linear models trained using just empirical risk minimization (ERM) and weight decay (so as to ensure a unique solution) have non-trivial correlation with the high variance feature in the data.
A key aspect of our analysis is that we can show a clear separation in terms of feature learning between Mixup/label smoothing and ERM in the setting of just linear models and linearly separable data distributions, whereas previous feature learning analyses of Mixup/similar (Chidambaram et al., 2023; Zou et al., 2023) have relied on data distributions that are explicitly constructed to not be learnable by linear models and have an asymptotically large number of classes. As a result, our analysis is both simpler than existing analyses and shows that nonlinear training dynamics are not necessary to separate Mixup/label smoothing and ERM with respect to feature learning.
We also verify that our theoretical results extend to practice across several experimental settings. In Section 4.1, we show that for synthetic data following our theoretical assumptions, logistic regression trained with Mixup or label smoothing learns a weight vector that is effectively uncorrelated with the high variance feature in the data. We also show that this behavior extends beyond synthetic data in Section 4.2, where we train logistic regression models on binary classification versions of CIFAR-10 and CIFAR-100 modified to have adversarial pixels in the training data that completely identify each class. In this case, the Mixup and label smoothing logistic regression models do not generalize at all due to having learned only the adversarial pixels in the training data.
Finally, we show in Section 4.3 that these observations still apply even when we consider ResNets trained on the full versions of CIFAR-10 and CIFAR-100; for these experiments we follow the experimental setup introduced by Chidambaram et al. (2023) in which the training data is modified to have extra low variance features via additional channels. Our results provide a contrasting perspective to that of Chidambaram et al. (2023) - Mixup (and label smoothing) can end up learning only the low variance features in the training data, whereas ERM can avoid this problem.
@@ 1.2 Related Work
Very recently, Oh & Yun (2023) studied Mixup in a similar context to our work (binary classification with linear models), and showed that Mixup can significantly improve the sample complexity required to learn the optimal classifier when compared to ERM. Also recently, and most closely related with our results in this paper, Mixup has been studied from a feature learning perspective, with Chidambaram et al. (2023) and Zou et al. (2023) both showing that Mixup training can learn multiple features in a generative, multi-view (Allen-Zhu & Li, 2021) data model despite ERM failing to do so. We provide a more detailed comparison to these works when we discuss our main results in Section 3.
@@ 2 Preliminaries
Notation. Given $n\in\mathbb{N}$, we use $[n]$ to denote the set $\{1,2,...,n\}$. For a vector $x\in\mathbb{R}^{d}$ and a subset $S\subset[d]$, we use $x_{S}\in\mathbb{R}^{\absolutevalue{S}}$ to denote the restriction of $x$ to only those indices in $S$, and also use $x_{i}$ to denote the $i^{\text{th}}$ coordinate of $x$. The same notation also applies to matrices; i.e. for a matrix $T$ we use $T_{S}$ to denote the restriction of the rows of $T$ to only those dimensions in $S$. Additionally, we use $\succ$ to denote the partial order over positive definite matrices and $\mathrm{Id}$ to denote the identity matrix. We use $\norm{\cdot}$ to indicate the Euclidean norm on $\mathbb{R}^{d}$. For two sets of points $A,B\subset\mathbb{R}^{d}$ we use $d(A,B):=\min_{x\in A,y\in B}\|x-y\|$ to denote the distance between $A$ and $B$. For a function $g:\mathbb{R}^{n}\to\mathbb{R}^{m}$ we use $g^{i}(x)$ to denote the $i^{\text{th}}$ coordinate function of $g$. For a probability distribution $\pi$ we use $\mathrm{supp}(\pi)$ to denote its support. Additionally, if $\pi$ corresponds to the joint distribution of two random variables $X$ and $Y$ (i.e. data and label), we use $\pi_{X}$ and $\pi_{Y}$ to denote the respective marginals, and $\pi_{X\mid Y=y}$ and $\pi_{Y\mid X=x}$ to denote the conditional distributions.
Our theory will focus on binary classification problems in $\mathbb{R}^{d}$ in which a subset of the input dimensions correspond to a low variance feature, and the complementary dimensions correspond to a higher variance but more separable feature. This setup is intended to mimic standard explorations of robustness of image classification models. For example, if each class in the data has a small set of constant, identifying pixels in every training point (either introduced adversarially or as an artifact of the domain being considered), does a model trained on this data learn to predict using only the identifying pixels or the whole image?
The goal in such cases is typically to also learn the features present in the non-identifying pixels (i.e. the higher variance features). We will show, however, that doing any kind of label augmentation (label smoothing or Mixup) in this low variance/high variance feature setup will lead to learning a model that has only learned the low variance feature, while standard training will lead to having non-trivial correlation with the high variance feature.
It is also natural to ask whether this phenomenon is purely a result of the fact that augmented labels (i.e. not 0 or 1) prevent the learned model weights from going to infinity, which has been pointed out to be crucial in previous existing analyses of Mixup (Chidambaram et al., 2023). We answer this question in the negative by also showing that training using ERM with weight decay still learns the high variance feature in the data, so the variance minimization effect is more than just a weight norm constraint.
2.1 Data Model
We begin with the generative model of the data that we will work with. For simplicity, we consider the low variance feature to have zero variance conditional on the label $y$ it is not difficult to extend our results to the case where the variance of this feature is simply much less than that of the high variance feature, but considering zero conditional variance yields cleaner proofs. We also consider balanced classes to further simplify the proofs, but our results will hold so long as each class appears with some non-zero probability.
@@ 2.1 Data Model
[Data Distribution] Let $\mathcal{L}\subset[d]$ and $\mathcal{H}=\mathcal{L}^{c}$ be the sets denoting the low variance feature and high variance feature in the data respectively. We define $\pi$ to be any distribution on $\mathbb{R}^{d}\times\{-1,1\}$ with $E_{x\sim\pi_{X}}[x]=0$ and $\pi_{Y}(y)=1/2$ for both $y\in\{-1,1\}$, along with the following properties:
  1. 1.
    (Zero Conditional Variance in $\mathcal{L}$) For $\pi_{X}$-a.e. $x\in\mathrm{supp}(\pi_{X\mid Y=y})$, we have that there exists a $v_{y}^{*}\in\mathbb{R}^{\absolutevalue{\mathcal{L}}}$ such that $x_{\mathcal{L}}=v_{y}^{*}$ with $\norm{v_{1}^{*}-v_{-1}^{*}}>0$.
  2. 2.
    (Non-zero Conditional Variance in $\mathcal{H}$) Let $\mu_{y}=\mathbb{E}[X\mid Y=y]$ and $\Sigma_{y}=\mathbb{E}[(x-\mu_{y})(x-\mu_{y})^{\top}\mid Y=y]$, then $\Sigma_{y,\mathcal{H}}$ (the restriction of $\Sigma_{y}$ to $\mathcal{H}$) is positive definite.
  3. 3.
    (More Separable High Variance Feature) Let $A_{y}=\{x_{\mathcal{H}}:x\in\mathrm{supp}(\pi_{X\mid Y=y})\}$. Then $A_{1},A_{-1}$ are compact, convex, and disjoint and satisfy $d(A_{1},A_{-1})>\norm{v_{1}^{*}-v_{-1}^{*}}$.
Since Definition 2.1 corresponds to a linearly separable data distribution, we can restrict our attention to linear models of the form $g_{w}(x)=\sigma(w^{\top}x)$, where $\sigma(\cdot)$ is the sigmoid function. We can then define the binary cross-entropy with weight decay as:
$\ell_{\lambda}(w)$ $=\mathbb{E}_{(x,y)\sim\pi}\left[-\log g_{w}(yx)\right]+\frac{\lambda}{2}\norm{w}^{2}$ (2.1)
With $\ell_{0}(w)$ thereby indicating the regular binary cross-entropy without weight decay. In what follows, we will work with the population form of the loss as defined in Equation (2.1), since for $\lambda>0$ we have that $\ell_{\lambda}(w)$ is strictly convex and therefore we may simply analyze its unique minimizer as opposed to worrying about training dynamics.
Although $\ell_{0}(w)$ will not have a unique minimizer, our results will still apply to the common case where $\ell_{0}(w)$ is minimized by scaling the max-margin solution, which we define below.
Definition 2.2.
[Max-Margin Solution] The max-margin solution $w^{*}$ with respect to $\pi$ is defined to be:
@@
2.2 Label Smoothing
Label smoothing as introduced by Szegedy et al. (2015) is intended to work with multi-class classification problems in which the data distribution $\pi$ is typically defined on $\mathbb{R}^{d}\times[k]$. In this case, we work with models $g:\mathbb{R}^{d}\to\Delta^{k}$ (the probability simplex in $\mathbb{R}^{k}$), and the cross-entropy loss is simply $\ell(g)=\mathbb{E}_{(x,y)\sim\pi}[-\log g^{y}(x)]$.
The label-smoothed cross-entropy is then obtained by replacing the one-hot encoded label $y$ with a mixture of the original label and the uniform distribution over $[k]$. Namely, the label-smoothed loss with mixing hyperparameter $\alpha\in[0,1]$ is defined to be:
$\ell_{\mathrm{LS},\alpha}(g)=-\mathbb{E}_{(x,y)\sim\pi}\bigg{[}$ $(1-\alpha)\log g^{y}(x)+\frac{\alpha}{k}\sum_{i=1}^{k}\log g^{i}(x)\bigg{]}$ (2.3)
For the purposes of our theory, however, we will be working with models of the form $g_{w}(x)=\sigma(w^{\top}x)$ and labels $y\in\{-1,1\}$, so we will consider the binary label-smoothed loss defined as:
$\ell_{\mathrm{LS},\alpha}(w)=-\mathbb{E}_{(x,y)\sim\pi}\bigg{[}$ $\left(1-\frac{\alpha}{2}\right)\log g_{w}(yx)+\frac{\alpha}{2}\log g_{w}(-yx)\bigg{]}$ (2.4)
Note that similar to Equation (2.1) the latter is defined as a function of the weight vector $w$ instead of the full model $g$.
2.3 Mixup
Similar to label smoothing, Mixup (Zhang et al., 2018) was also introduced in the context of multi-class classification. In addition to augmenting the one-hot encoded labels $y$, Mixup also augments the input data $x$ by combining both with another data point $(x^{\prime},y^{\prime})$ via a random convex combination. Both of these transformations are done according to a mixing distribution $\mathcal{D}_{\lambda}$ (which is a hyperparameter) whose support is contained in $[0,1]$. To simplify notation, we will use $x_{1:n}$ and $y_{1:n}$ to denote multiple data points $x_{1},...,x_{n}$ and their corresponding labels $y_{1},...,y_{n}$. Additionally, in the multi-class setting with $g:\mathbb{R}^{d}\to\Delta^{k}$, we also define:
$h_{\lambda}(g,x_{1:2},y_{1:2})$ $=\lambda\log g^{y_{1}}(z_{\lambda})+(1-\lambda)\log g^{y_{2}}(z_{\lambda})$ (2.5)
$\text{where}\quad z_{\lambda}$ $=\lambda x_{1}+(1-\lambda)x_{2}$ (2.6)
We may then define the Mixup cross-entropy:
$\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(g)=\mathbb{E}_{(x_{1:2},y_{1:2})\sim\pi,\lambda\sim\mathcal{D}_{\lambda}}\left[-h_{\lambda}(g,x_{1:2},y_{1:2})\right]$ (2.7)
Where $(x_{1:2},y_{1:2})\sim\pi$ is shorthand for $(x_{1},y_{1}),(x_{2},y_{2})\sim\pi$. As before, since our theory focuses on binary classification and linear models $g_{w}$, we may identically define the binary version $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ of Equation (2.7) except for redefining $h_{\lambda}$ to be:
$h_{\lambda}(w,x_{1:2},y_{1:2})$ $=\lambda\log g_{w}(y_{1}z_{\lambda})+(1-\lambda)\log g_{w}(y_{2}z_{\lambda})$ (2.8)
Again, observe that we have redefined $h_{\lambda}$ to directly be a function of $w$ instead of $g$.
@@ 3 Main Theoretical Results
Having established the necessary background, we now state our main results. Full proofs of all results can be found in Appendix A. We first show that the unique minimizer $w^{*}$ of the standard risk with weight decay (as defined in Equation (2.1)) has non-trivial correlation with the high variance dimensions of the input $x$.

### v2

@@ Abstract
Data augmentation has been pivotal in successfully training deep learning models on classification tasks over the past decade. An important subclass of data augmentation techniques - which includes both label smoothing and Mixup - involves modifying not only the input data but also the input label during model training. In this work, we analyze the role played by the label augmentation aspect of such methods. We first prove that linear models on binary classification data trained with label augmentation learn only the minimum variance features in the data, while standard training (which includes weight decay) can learn higher variance features. We then use our techniques to show that even for nonlinear models and general data distributions, the label smoothing and Mixup losses are lower bounded by a function of the model output variance. An important consequence of our results is negative: label smoothing and Mixup can be less robust to spurious correlations in the data. We verify that our theory reflects practice via experiments on image classification benchmarks modified to have spurious correlations.
@@ 1 Introduction
Perhaps the two most widely applied data augmentation methods in this subcategory are label smoothing (Szegedy et al., 2015) and Mixup (Zhang et al., 2018). Label smoothing replaces the one-hot encoded labels in the training data with smoothed out labels that assign non-zero probability to every possible class (see Section 2 for a formal definition). Mixup similarly smooths out the training labels, but does so via introducing random convex combinations of data points and their labels. As a result, Mixup modifies not only the training labels but also the training inputs. The general principles of label smoothing and Mixup have been extended to several variants, such as Structural Label Smoothing (Li et al., 2020), Adaptive Label Smoothing (Wang et al., 2021), Manifold Mixup (Verma et al., 2019), CutMix (Yun et al., 2019), PuzzleMix (Kim et al., 2020), SaliencyMix (Uddin et al., 2020), AutoMix (Liu et al., 2021), and Noisy Feature Mixup (Lim et al., 2022).
@@
1.1 Main Contributions
Our main contributions are three-fold:
  1. 1.
    We introduce a simple class of data distributions in which there exists a low variance feature and a high variance feature, and show that optimizing the label smoothing and Mixup losses for a linear model for such distributions leads to learning only the low variance feature. On the other hand, if the high variance feature satisfies a greater separability condition relative to the low variance feature, optimizing the empirical risk with weight decay successfully learns both features in the data. Our label smoothing and Mixup results are stated under general conditions on both the data and the respective loss hyperparameters; our Mixup results apply to any mixing distribution that does not coincide with ERM, which greatly generalizes the settings considered by prior work on Mixup feature learning (Chidambaram et al., 2023; Zou et al., 2023).
  2. 2.
    We show that our techniques extend to the analysis of general model classes, and prove lower bounds for the label smoothing and Mixup losses in terms of the model output variance. These results suggest that optimizing these losses necessarily decreases model output variance, which is not necessarily the case for training with weight decay.
  3. 3.
    We empirically verify our theoretical insights by considering image classification benchmarks modified to (1) have adversarial spurious correlations in the training data and (2) more natural spurious correlations in the form of class-correlated backgrounds. In both cases, we observe that label smoothing and Mixup fail to generalize for the full range of hyperparameters used in practice, while ERM with weight decay does generalize.
@@ 1.2 Related Work
Recently, Oh & Yun (2023) studied Mixup in a similar context to our work (binary classification with linear models), and showed that Mixup can significantly improve the sample complexity required to learn the optimal classifier when compared to ERM. Most closely related with our results in this paper, Mixup has been studied from a feature learning perspective, with Chidambaram et al. (2023) and Zou et al. (2023) both showing that Mixup training can learn multiple features in a generative, multi-view (Allen-Zhu & Li, 2021) data model despite ERM failing to do so. We provide a more detailed comparison to these works when we discuss our main results in Section 3.
@@ 2 Preliminaries
[…]
We consider the $k$-class classification setting, in which there is a ground-truth data distribution $\pi$ on $\mathbb{R}^{d}\times[k]$ and our goal is to model the conditional distribution $\pi_{Y\mid X}$ using a learned function $g$. In our main theoretical results, we will pay particular attention to the case where $k=2$ and $g$ is a logistic regression model (although we generalize a weak version of these observations to $k$ classes as well), as in this setting we can get a clear handle on the features in the data learned by an optimal model with respect to a particular loss. For this case, we will assume that $\pi$ is supported on $\mathbb{R}^{d}\times\{\pm 1\}$ and that $g$ is parameterized by a weight vector $w$, i.e. $g_{w}(x)=\sigma(w^{\top}x)$, where $\sigma$ is the sigmoid function. Throughout this work, we will consider the following three families of losses.
Standard cross-entropy with optional weight decay. The canonical cross-entropy (or negative log-likelihood) objective in the $k$-class setting is defined as:
$\ell(g)=\mathbb{E}_{(X,Y)\sim\pi}[-\log g^{Y}(X)].$ (2.1)
Here we have not specified a weight decay term, since we have placed no constraints on the structure of $g$. On the other hand, for linear binary classification, we can define the binary cross-entropy with optional weight decay as (recalling that $Y\in\{\pm 1\}$)
$\ell_{\beta}(w)$ $=\mathbb{E}_{(X,Y)\sim\pi}\left[-\log g_{w}(YX)\right]+\frac{\beta}{2}\norm{w}^{2},$ (2.2)
where we have defined the loss in terms of the parameter vector $w$. When $\beta>0$, (2.2) has a unique minimizer and we will directly analyze its properties. On the other hand, when $\beta=0$ (no weight decay), this is no longer the case - but our results will still apply to the common case of minimizing (2.2) by scaling the max-margin solution, which we define below.
@@
[Max-Margin Solution] The max-margin solution $w^{*}$ with respect to $\pi$ is defined as:
@@
Label smoothing. The cross-entropy as defined in (2.1) treats the reference distribution that we compare $g(X)$ to as a point mass on the class $Y$. On the other hand, the label-smoothed cross-entropy is obtained by instead treating the reference distribution as a mixture of a point mass on $Y$ and the uniform distribution over $[k]$. Namely, the label-smoothed loss with mixing hyperparameter $\alpha\in[0,1]$ is defined to be:
$\ell_{\mathrm{LS},\alpha}(g)=-\mathbb{E}_{(X,Y)\sim\pi}\bigg{[}(1-\alpha)\log g^{Y}(X)+\frac{\alpha}{k}\sum_{i=1}^{k}\log g^{i}(X)\bigg{]}.$ (2.4)
And the corresponding binary version is (once again, $Y\in\{\pm 1\}$):
$\ell_{\mathrm{LS},\alpha}(w)=-\mathbb{E}_{(X,Y)\sim\pi}\bigg{[}$ $\left(1-\frac{\alpha}{2}\right)\log g_{w}(YX)+\frac{\alpha}{2}\log g_{w}(-YX)\bigg{]}.$ (2.5)
Mixup. Similar to label smoothing, Mixup also augments the reference distribution from being a point mass on $Y$ to being a mixture. However, Mixup also augments the input data as well. In particular, Mixup considers convex combinations of two pairs of points $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$, with a mixing weight sampled from a distribution $\mathcal{D}_{\lambda}$ (which is a hyperparameter) whose support is contained in $[0,1]$. To simplify notation, we will use $X_{1:n}$ and $Y_{1:n}$ to denote multiple inputs $X_{1},...,X_{n}$ and their corresponding labels $Y_{1},...,Y_{n}$, and additionally introduce a function $h$ defined as:
$h(\lambda,g,X_{1:2},Y_{1:2})$ $=\lambda\log g^{Y_{1}}(Z_{\lambda})+(1-\lambda)\log g^{Y_{2}}(Z_{\lambda})$ (2.6)
$\text{where}\quad Z_{\lambda}$ $=\lambda X_{1}+(1-\lambda)X_{2}.$ (2.7)
After which we can define the Mixup cross-entropy as:
$\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(g)=\mathbb{E}_{(X_{1:2},Y_{1:2})\sim\pi,\lambda\sim\mathcal{D}_{\lambda}}\left[-h(\lambda,g,X_{1:2},Y_{1:2})\right].$ (2.8)
The corresponding binary version $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ of (2.8) is identical except for redefining $h$ to be:
$h(\lambda,w,x_{1:2},y_{1:2})$ $=\lambda\log g_{w}(Y_{1}Z_{\lambda})+(1-\lambda)\log g_{w}(Y_{2}Z_{\lambda}).$ (2.9)
@@ 3 Main Theoretical Results
All omitted proofs in this section can be found in Section A of the Appendix.
3.1 Linear Binary Classification
We begin first with the binary setting, in which we will consider a data model where a subset of the input dimensions correspond to a low variance feature and the complementary dimensions correspond to a high variance feature that is more separable. This setup is intended to be a simplistic model of spurious correlations that appear in image data; one could think of the low variance feature as a fixed background while the high variance feature corresponds to the actual object category of interest.
The goal in such cases is to ensure that we do not only learn the low variance feature, so that generalization is possible even under a shift of this feature (for example, recognizing the same object on different backgrounds). We will show, however, that doing any kind of label augmentation (label smoothing or Mixup) in this low variance/high variance feature setup will lead to a model that has only learned the low variance feature, whereas minimizing the binary cross-entropy with non-zero weight decay requires learning the high variance feature due to its greater separability. The overarching data model for this subsection is as follows.
Definition 3.1.
[Binary Data Distribution] We consider $\pi$ to be a distribution supported on $K\times\{\pm 1\}$ where $K$ is a compact subset of $\mathbb{R}^{d}$. We assume that $\pi$ is nondegenerate in that it satisfies $\pi_{Y}(y)>0$ for each $y\in\{\pm 1\}$ and that $\Sigma_{X}$ is positive definite, and we also assume $\mathbb{E}[X]=0$. Additionally, we designate a subset $\mathcal{L}\subseteq[d]$ that we refer to as the low variance feature in the data, and we refer to the complement $\mathcal{H}=\mathcal{L}^{c}$ as the high variance feature.
For convenience, for a vector $v\in\mathbb{R}^{d}$, we will use $v\in\mathcal{L}$ to mean that only the $\mathcal{L}$ dimensions of $v$ are non-zero. Definition 3.1 does not actually make any strong assumptions on $\pi$ yet; this is because the weight decay and label augmentation results have different, non-conflicting assumptions that we codify separately. For the former, we need the data to be more separable in the $\mathcal{H}$ dimensions than in the $\mathcal{L}$ dimensions, in the following sense.
Assumption 3.2.
We assume that for every unit vector $u^{*}\in\mathcal{L}$, there exists a unit vector $v^{*}\in\mathcal{H}$ and $y\left\langle v^{*},x\right\rangle>y\left\langle u^{*},x\right\rangle$ for $\pi$-a.e. $(x,y)$.
Assumption 3.2 is strong, but it actually does not imply linear separability, only that the $\mathcal{H}$ dimensions are in a sense better than the $\mathcal{L}$ dimensions for classification. For a simple example to provide intuition for the assumption, one can consider a 2-D distribution in which the $x$ coordinates for both classes are concentrated close to 0, but the $y$ coordinates are either large negative or large positive values depending on the class.
Our first result shows that for $\ell_{\beta}(w)$ as defined in (2.2), the minimizer $w^{*}$ has a large correlation with the dimensions in $\mathcal{H}$. This is of course intuitive given Assumption 3.2, but it is not immediate because the weight decay penalty in (2.2) encourages distributing norm across all of the dimensions of $w$ (e.g. $\sum w_{i}$ is maximized with respect to the constraint $\norm{w}=1$ by considering $w_{i}=1/\sqrt{d}$).